OpenStudy (seattle12345):

Can someone please help with this: Find all real solutions of the equation: 2(c-2)^2-7=9

OpenStudy (anonymous):

is this your question?? \[\Large 2(c-2)^2-7=9\]

OpenStudy (seattle12345):

yes

OpenStudy (anonymous):

ok using \[\Large (a-b)^2=a^2+b^2-2ab\] expand (c-2)^2 as \[\Large 2(c^2+4-4c)-7=9\] now simplify\[\Large \implies (2c^2+8-8c)-7=9\] \[\Large \implies 2c^2-8c-8=0\] it is Quadratic Equation

OpenStudy (anonymous):

@seattle12345 can you do this now ?

OpenStudy (anonymous):

@seattle12345 waiting for your reply!

OpenStudy (seattle12345):

sorry I was working on it. so we need two factors of -16 whose sum is -8 right?

OpenStudy (anonymous):

ok i think you should use Quadratic Equation these cannot be factored. are you familiar with quadratic Equation ?

OpenStudy (seattle12345):

no we have not gone over that yet

OpenStudy (anonymous):

completing squares??? you familiar with completing squares?

OpenStudy (seattle12345):

haven't covered that either

OpenStudy (anonymous):

ok original equation is \[\Large 2(c-2)^2-7=9\] add seven to both sides \[\Large \implies 2(c-2)^2-7+7=9+7\] \[\Large \implies 2(c-2)^2=16\] divide both sides by 2 \[\Large \implies \frac{2(c-2)^2}{2}=\frac{16}{2}\] \[\Large \implies (c-2)^2=8\] taking square root of both sides \[\Large \implies \sqrt{(c-2)^2}=\sqrt{8}\] \[\Large \implies (c-2)=\pm \sqrt{8}\] can you do this now ??

OpenStudy (anonymous):

@seattle12345

OpenStudy (seattle12345):

yes! thank you!

OpenStudy (anonymous):

you are welcome:)